Find four consecutive integers such that the sum of twice the first and three times the third is 126.

Answer by macston(5194)

F=first integer; F+2=Third integer
2F+3(F+2)=126
2F+3F+6=126 Subtract 6 from each side.
5F=120 Divide each side by 5
F=24 ANSWER 1: The first integer is 24
F+1=25 ANSWER 2: The second integer is 25
F+2=26 ANSWER 3: Third integer is 26
F+3=27 ANSWER 4: Fourth integer is 27
CHECK:
Twice first + 3 times Third = 126
2(24)+3(26)=126
48+78=126
126=126